OpenStudy (anonymous):

The ratio of the areas of two circles is 9 to 16. If the circumference of the larger circle is 20 cm, what is teh circumference of the smaller circle?

OpenStudy (anonymous):

Are you familiar with cross multiplying? I belive you go 9/16 times (x/20cm) From there, working it out it becomes (20x9)/16 and that gave me 45/4cm or 11.25

OpenStudy (anonymous):

so answer is 11.25 cm?

OpenStudy (anonymous):

yeah thats what I got, still though do you know how to cross multiply?

OpenStudy (anonymous):

yes i noe, y?

OpenStudy (anonymous):

makes life much easier, just making sure because thats usuallly all you do when you have a ratio problem like this

OpenStudy (anonymous):

ok, but i just realized that the ratio given about the two circles is their area, not circumference!

OpenStudy (anonymous):

area=pi*r*r circumference is pi*2*r ratio of area of circle 1 to circle 2 is 9:16. 9/16=area1/area2=pi*r1*r1/pi*r2*r2=r1*r1/r2*r2=3*3/4*4 so, ratio of r is 3:4

OpenStudy (anonymous):

is teh answer 11.25 still correct then?

OpenStudy (anonymous):

lyam, sdo wat's the final answer?

OpenStudy (anonymous):

nope 3/4=X/20 4X=60 X=60/4=15 I think

OpenStudy (anonymous):

sorry i type slow